# Berry phases and connection matrices defined on homogeneous spaces attached to Siegel-Jacobi groups

### v1

@@ 2. Preparation
Note that (1) is different from (2.25b) and (2.49). The matrix of the balanced metric $h=h(\varsigma)$, $\varsigma:=(z,w)\in\mathbb{C}\times{\mathcal{D}}_{1}$ associated to the Kähler two-form (2.11) is
@@
Introducing the second partial Cayley transform (2.24a) into the Kähler two-form (2.11) on ${\mathcal{D}}^{J}_{1}$, we get the symplectic two-form (1) on the Siegel-Jacobi upper half-plane $(v,\eta)$, $\operatorname{Im}v>0$
@@
The Berndt–Kähler’s two-form (symplectic two-form) invariant to the action of $G^{J}(\mathbb{R})_{0}$ =SL(2,)absentleft-normal-factor-semidirect-productnormal-$=\rm{SL}(2,\mathbb{R})\ltimes\mathbb{C}$, is (2.31a) ((2.31c))
@@
and finally we regain (1).
The matrix corresponding to the balanced metric (1) associated with the Kähler  two -form (2.31a) reads
@@
The metric matrix associated with (1) is
@@
Formula (2.41) is presented by Berndt as “communicated to the author by Kähler”. Also in [39, p 8] is given our equation (5.21a) in first reference [27], while our present equation (1) corrects two printing errors in Berndt’s paper.
@@
The metric (8) in [61] differs from the metric (1) by a factor of two, since the Hermitian metric used by Kähler  is $\operatorname{d}s^{2}=2g_{i\bar{j}}\operatorname{d}z_{i}\otimes\operatorname{d}\bar{z}_{j}$. If in (2.42) we take $k/2\rightarrow k$, we have
@@
and we get (2.31a). Relation (2.31a) has been obtained by Berndt [38, p 30], where the denominator of the first term is misprinted as $v-\bar{v}$ (or $\tau-\bar{\tau}$ in our notations). Equation (1) appears also in [38, p 30] and [40, p 62].
We denote in (2.42) $(\tau,z)$ with $(v,u)$ as in [15, (9.16)]. Indeed, we make successively the transformations: the partial Cayley transform, $\Phi:~{}(w,z)\rightarrow(v,u)$ (1), a holomorphic transform, and we get (2.44a), then we apply the $FC_{1}$ transform $(v,u)\rightarrow(v,\eta)$ (2.34), a non-holomorphic transform, to obtain (2.44b), and finally we make the symplectic transform $(w,z)\rightarrow(x,y,q,p)$ with the result (2.44c)
@@
The metric (1) corresponds to the Kähler potential (2.44a)
@@
Yang calculated in [92] the metric on ${\mathcal{X}}^{J}_{n}$, invariant to the action of $G^{J}_{n}(\mathbb{R})_{0}$. The equivalence of the metric of Yang with the metric obtained via CS on ${\mathcal{D}}^{J}_{n}$ and then transported to ${\mathcal{X}}^{J}_{n}$ via the partial Cayley transform $(v,u)\rightarrow(v,\eta)$ (1) is underlined in [19]. In particular, the metric (5.21b) in the first reference [27] appears in [92, p 99] for the particular values $c_{1}=1$, $c_{2}=4$. See also [91, 93, 94].
@@ 3.1. Balanced metric
The Kähler two-form $\omega_{{\mathcal{D}}^{J}_{1}}(w,\eta)$ given by (2.14a) ($\omega(v,\eta)$, (1)) can be obtained from the Kähler  potentials (3.7a) (respectively (3.7b)) using a formula of the type (3.2b)
@@ 3.3. Linear Hamiltonian in the generators of the Jacobi group $G^{J}_{1}(\mathbb{R})$
We consider formula (1) of the transformation $w\rightarrow v$ and we write this change of variables as
@@
We apply relation (5.7) to the change of variables $v\rightarrow w$ (1). We have
@@ 3.5. Connection matrices on ${\mathcal{D}}^{J}_{1}$ and ${\mathcal{X}}^{J}_{1}$
The non-zero Christoffel’s symbols corresponding to the Riemannian metric $\operatorname{d}\!s^{2}_{{\mathcal{X}}^{J}_{1}}\!(\!x\!,y\!,\!q,\!p\!)$ (1) on the Siegel-Jacobi upper half-plane are [29, (73)]
@@
which is true because of of (1), (3.35).
@@ 4.1. Definitions
${\mathrm{vec}}(A)^{t}=[a_{11},a_{21},\dots,a_{1n},a_{21},\dots,a_{2n},a_{31},\dots,a_{nn}]\in M(1,n^{2},\mathbb{C}),$
@@
${\mathrm{vech}}(A)^{t}=[a_{11},a_{12},\dots,a_{1n},a_{22},\dots,a_{2n},a_{33},\dots,a_{nn}]\in M(1,N_{1}),~{}N_{1}=\frac{n(n+1)}{2}.$
@@
(4.4) $\Omega=\sum_{I=1}^{{\mathrm{n}}}\operatorname{d}Q^{I}\wedge\operatorname{d}P^{I}.$
@@
If in formula (3) of the Kähler-two form on ${\mathcal{D}}^{J}_{n}$ we make the partial the Cayley transform $(v,u)\rightarrow(v,\eta)=(x+\operatorname{i}y,q+\operatorname{i}p)$ (3), where $v\in MS(n,\mathbb{C})$, we get the Kähler  two-form $\omega_{{\mathcal{X}}^{J}_{n}}$
@@
We arrange the elements of the symmetric matrices $x,y^{-1}$ in $\rm{vech}(x),\rm{vech}(y^{-1})$ s.t. there are $i$ elements on every row, $i=1$ on the last row, and $i=n$ on the first row. We use the notation
@@
$\operatorname{tr}(G^{t}D\wedge\bar{G})$ $=\frac{\operatorname{i}}{2}\operatorname{tr}\{[\operatorname{d}p(x+\operatorname{i}y)+\operatorname{d}q]y^{-1}\wedge[(x-\operatorname{i}y)\operatorname{d}p^{t}+\operatorname{d}q^{t}]$
$=\frac{\operatorname{i}}{2}\operatorname{tr}\{\operatorname{d}p(x\!+\!\operatorname{i}y)y^{-1}\wedge[\!(x\!-\!\operatorname{i}y)\operatorname{d}p^{t}\!+\!\operatorname{d}q^{t}\!]\!+\!\operatorname{d}qy^{-1}\wedge[(\!x-\!\operatorname{i}y\!)\operatorname{d}p^{t}\!+\!\operatorname{d}q^{t}\!]$
@@
$+\operatorname{d}q[(y^{-1}x-\operatorname{i})\wedge\operatorname{d}p^{t}+q\wedge y^{-1}\operatorname{d}q^{t}]\}$
$=\frac{\operatorname{i}}{2}\operatorname{tr}\{\operatorname{d}p(xy^{-1}x+yy^{-1}y)\operatorname{d}p^{t}+\operatorname{d}pxy^{-1}\wedge\operatorname{d}q^{t}$
@@
Lemma 3, extracted from [30, Lemma 1], is a particular case of Lemma 2. To proof (2.25b) we introduce the relations (2.24a) into (1).
@@ 5.1.1. Connections on real manifolds
For (5.11), equation (5.14) with (5.2), (1) becomes
@@
In [60, §4] the affine connection $D$ is denoted $\nabla$, $D_{X}$ defined at (5.2) is denoted $\nabla_{X}$ and (1) becomes
@@ 5.2. Examples
Formula (5.61a) of the Berry connection on the complex Grassmann manifold $G_{n}(\mathbb{C}^{n+m})$ corresponding to the scalar product (1) was obtained in [35, (5.17)]

### v2

# Berry phases and connection matrices defined on homogeneous spaces attached to Jacobi groups

@@ 2. Preparation
Note that (2.20) is different from (2.25b) and (2.49). The matrix of the balanced metric $h=h(\varsigma)$, $\varsigma:=(z,w)\in\mathbb{C}\times{\mathcal{D}}_{1}$ associated to the Kähler two-form (2.11) is
@@
Introducing the second partial Cayley transform (2.24a) into the Kähler two-form (2.11) on ${\mathcal{D}}^{J}_{1}$, we get the symplectic two-form (2.25) on the Siegel-Jacobi upper half-plane $(v,\eta)$, $\operatorname{Im}v>0$
@@
The Berndt–Kähler’s two-form (symplectic two-form) invariant to the action of $G^{J}(\mathbb{R})_{0}$ $=\rm{SL}(2,\mathbb{R})\ltimes\mathbb{C}$, is (2.31a) ((2.31c))
@@
and finally we regain (2.25).
The matrix corresponding to the balanced metric (2.46) associated with the Kähler  two -form (2.31a) reads
@@
The metric matrix associated with (2.39) is
@@
Formula (2.41) is presented by Berndt as “communicated to the author by Kähler”. Also in [39, p 8] is given our equation (5.21a) in first reference [27], while our present equation (2.39) corrects two printing errors in Berndt’s paper.
@@
The metric (8) in [61] differs from the metric (2.39) by a factor of two, since the Hermitian metric used by Kähler  is $\operatorname{d}s^{2}=2g_{i\bar{j}}\operatorname{d}z_{i}\otimes\operatorname{d}\bar{z}_{j}$. If in (2.42) we take $k/2\rightarrow k$, we have
@@
and we get (2.31a). Relation (2.31a) has been obtained by Berndt [38, p 30], where the denominator of the first term is misprinted as $v-\bar{v}$ (or $\tau-\bar{\tau}$ in our notations). Equation (2.39) appears also in [38, p 30] and [40, p 62].
We denote in (2.42) $(\tau,z)$ with $(v,u)$ as in [15, (9.16)]. Indeed, we make successively the transformations: the partial Cayley transform, $\Phi:~{}(w,z)\rightarrow(v,u)$ (2.28), a holomorphic transform, and we get (2.44a), then we apply the $FC_{1}$ transform $(v,u)\rightarrow(v,\eta)$ (2.34), a non-holomorphic transform, to obtain (2.44b), and finally we make the symplectic transform $(w,z)\rightarrow(x,y,q,p)$ with the result (2.44c)
@@
The metric (2.46) corresponds to the Kähler potential (2.44a)
@@
Yang calculated in [92] the metric on ${\mathcal{X}}^{J}_{n}$, invariant to the action of $G^{J}_{n}(\mathbb{R})_{0}$. The equivalence of the metric of Yang with the metric obtained via CS on ${\mathcal{D}}^{J}_{n}$ and then transported to ${\mathcal{X}}^{J}_{n}$ via the partial Cayley transform $(v,u)\rightarrow(v,\eta)$ (2.28) is underlined in [19]. In particular, the metric (5.21b) in the first reference [27] appears in [92, p 99] for the particular values $c_{1}=1$, $c_{2}=4$. See also [91, 93, 94].
@@ 3.1. Balanced metric
The Kähler two-form $\omega_{{\mathcal{D}}^{J}_{1}}(w,\eta)$ given by (2.14a) ($\omega(v,\eta)$, (2.25)) can be obtained from the Kähler  potentials (3.7a) (respectively (3.7b)) using a formula of the type (3.2b)
@@ 3.3. Linear Hamiltonian in the generators of the Jacobi group $G^{J}_{1}(\mathbb{R})$
We consider formula (2.24) of the transformation $w\rightarrow v$ and we write this change of variables as
@@
We apply relation (5.7) to the change of variables $v\rightarrow w$ (2.24). We have
@@ 3.5. Connection matrices on ${\mathcal{D}}^{J}_{1}$ and ${\mathcal{X}}^{J}_{1}$
The non-zero Christoffel’s symbols corresponding to the Riemannian metric $\operatorname{d}\!s^{2}_{{\mathcal{X}}^{J}_{1}}\!(\!x\!,y\!,\!q,\!p\!)$ (2.39) on the Siegel-Jacobi upper half-plane are [29, (73)]
@@
which is true because of of (2.28), (3.35).
@@ 4.1. Definitions
$({\mathrm{vec}}(A)^{t})^{t}=[a_{11},a_{12},\dots,a_{1n},a_{21},\dots,a_{2n},a_{31},\dots,a_{nn}]\in M(1,n^{2},\mathbb{C}),$
@@
$({\mathrm{vech}}(A)^{t})^{t}=[a_{11},a_{12},\dots,a_{1n},a_{22},\dots,a_{2n},a_{33},\dots,a_{nn}]\in M(1,N_{1}),~{}N_{1}=\frac{n(n+1)}{2}.$
@@
(4.4) $\Omega=\sum_{I=1}^{{\mathrm{\frac{n(n+3)}{2}}}}\operatorname{d}Q^{I}\wedge\operatorname{d}P^{I}.$
@@
If in formula (2.53) of the Kähler-two form on ${\mathcal{D}}^{J}_{n}$ we make the partial the Cayley transform $(v,u)\rightarrow(v,\eta)=(x+\operatorname{i}y,q+\operatorname{i}p)$ (2.54), where $v\in MS(n,\mathbb{C})$, we get the Kähler  two-form $\omega_{{\mathcal{X}}^{J}_{n}}$
@@
We arrange the elements of the symmetric matrices $x,y^{-1}$ in $((\rm{vech}(x))^{t})^{t},((\rm{vech}(y^{-1}))^{t})^{t}$ s.t. there are $i$ elements on every row, $i=1$ on the last row, and $i=n$ on the first row. We use the notation
@@
$\operatorname{tr}(G^{t}D\wedge\bar{G})$ $=\frac{\operatorname{i}}{2}\operatorname{tr}\{[\operatorname{d}p(x+\operatorname{i}y)+\operatorname{d}q]y^{-1}\wedge[(x-\operatorname{i}y)\operatorname{d}p^{t}+\operatorname{d}q^{t}]\}$
$=\frac{\operatorname{i}}{2}\operatorname{tr}\{\operatorname{d}p(x\!+\!\operatorname{i}y)y^{-1}\wedge[\!(x\!-\!\operatorname{i}y)\operatorname{d}p^{t}\!+\!\operatorname{d}q^{t}\!]\!+\!\operatorname{d}qy^{-1}\wedge[(\!x-\!\operatorname{i}y\!)\operatorname{d}p^{t}\!+\!\operatorname{d}q^{t}\!]\}$
@@
$+\operatorname{d}q[(y^{-1}x-\operatorname{i})\wedge\operatorname{d}p^{t}+\wedge y^{-1}\operatorname{d}q^{t}]\}$
$=\frac{\operatorname{i}}{2}\operatorname{tr}\{\operatorname{d}p(xy^{-1}x+yy^{-1}y)\wedge\operatorname{d}p^{t}+\operatorname{d}pxy^{-1}\wedge\operatorname{d}q^{t}$
@@
Lemma 3, extracted from [30, Lemma 1], is a particular case of Lemma 2. To proof (2.25b) we introduce the relations (2.24a) into (2.25).
@@ 5.1.1. Connections on real manifolds
For (5.11), equation (5.14) with (5.2), (5.1) becomes
@@
In [60, §4] the affine connection $D$ is denoted $\nabla$, $D_{X}$ defined at (5.2) is denoted $\nabla_{X}$ and (5.1) becomes
@@ 5.2. Examples
Formula (5.61a) of the Berry connection on the complex Grassmann manifold $G_{n}(\mathbb{C}^{n+m})$ corresponding to the scalar product (5.1) was obtained in [35, (5.17)]